How many ounces of a 10 percent solution and a 30 percent solution
mix to make 20 ounces of a 25 percent solution

Answer by stanbon(75887)

How many ounces of a 10 percent solution and a 30 percent solution
mix to make 20 ounces of a 25 percent solution
-------
Equation:
active + active = active
0.10x + 0.30(20-x) = 0.25*20
---------------------------
Multiply thru by 100 to get:
10x + 30*20 - 30x = 25*20
----
-20x = -5*20
x = 5 ounces (amt. of 10% solution needed)
----
20-x = 15 ounces (amt. of 30% solution needed)
=================
Cheers,
Stan H.
===================